In 1944 a stamp for first-class postage cost 3 cents. By the end of 2001, the cost of first-class postage will increase 35 cents. Assuming that the average annual rate of inflation was 5% for the period, is 35 cents a reasonable price in 2001? If not, what should the price be?
Could anyone please explain this to me? Thanks.

Answer by stanbon(48552)

In 1944 a stamp for first-class postage cost 3 cents. By the end of 2001, the cost of first-class postage will increase 35 cents. Assuming that the average annual rate of inflation was 5% for the period, is 35 cents a reasonable price in 2001? If not, what should the price be?
-------------------------
A(t) = A(o)(1.05)^t
'44 to '01 is 57 years
A(t) = 0.03(1.05)^57
A(t) = 0.03*16.14 = 0.48 or 48 cents
-----------
At an annual inflation rate of 5% per year for 57 years
a 3 cent stamp in 1944 would cost 48 cents in 2001.
So, it look like 35 cents is a bargain.
===============
Cheers,
Stan H.